Shadow:

Perform the indicated operation and state the domain. Let f(x) = 9x^-1 , g(x) = x^(1/4)

Shadow:

\[\frac{ f(x) }{ g(x) }\]

satellite73:

write one on top of the other

satellite73:

then subtract the exponents

Shadow:

\[\frac{ 9x ^{-1} }{ x ^{\frac{ 1 }{ 4 }} }\]

satellite73:

that is a good start

Shadow:

Couldn't I move the 9x^-1 down?

satellite73:

no

satellite73:

you can move bring the \(x^{-1}\) in the denominator as \(x\) but the 9 is not glued to it, it stays up top

Shadow:

oh you're right, mb

Shadow:

So would it be something like \[x ^{\frac{ 1 }{ 4 } - x^-1}\] Since 1/4 is the larger number

satellite73:

it would be \[\large \frac{9}{x^{\frac{5}{4}}}\] since \[\frac{1}{4}+1=\frac{5}{4}\]

Shadow:

Oh, right, I don't know why I had the x there. I was thinking that though

Shadow:

Is that as far as I can go?

Shadow:

I have to go, but I'll be back on to check the question in about an hour. Thank you for all your assistance. Have a test tomorrow.

satellite73:

yes that is all

Shadow:

Ok, thank you